OpenStudy (anonymous):

Patricia owned so many posters that she decided to sell 63 of them to her friends. After selling the posters, she still had 172 left. Write and solve an equation to find the number of posters p Patricia had originally P.S i know the answer just wanna see how many people can figure it out :D

OpenStudy (anonymous):

p = h + s, where h = 172, and s = 63 p = 172 + 63 = 235

OpenStudy (anonymous):

An easier way to solve this equation is do 63+172 then you know that P=235 before she sold the posters